please show me how to solve this problem: the perimeter of a rectangle is 50m. if the width were doubled and the length were increased by 25m, the perimeter would be 112m. what are the length and width of the rectangle?
Answer by scott8148(6628)

"the perimeter of a rectangle is 50m" ___ 2L + 2W = 50 (A)

"if the width were doubled and the length were increased by 25m, the perimeter would be 112m"
___ 2(L + 25) + 2(2W) = 112 ___ 2L + 50 + 4W = 112 (B)

subtracting equation A from equation B will eliminate L

find W, and then substitute back to find L